### v1

# Interval Regression: A Comparative Study with Proposed Models

Abstract
Regression models are essential for a wide range of real-world applications. However, in practice, target values are not always precisely known; instead, they may be represented as intervals of acceptable values. This challenge has led to the development of Interval Regression models. In this study, we provide a comprehensive review of existing Interval Regression models and introduce alternative models for comparative analysis. Experiments are conducted on both real-world and synthetic datasets to offer a broad perspective on model performance. The results demonstrate that no single model is universally optimal, highlighting the importance of selecting the most suitable model for each specific scenario.
keywords:
interval regression, survival analysis, supervised machine learning.
1 Introduction
Need for Interval Regression over Standard Regression
Regression models are widely used in real-world applications. However, due to some reasons, such as errors in measurement, the data collection procedures, or human factors, the target value in many scenarios is not a single value but rather an interval of acceptable values. For instance, in medical research [1, 2], the precise survival time of a patient is often uncertain, but it is known that the patient survives for at least a certain number of years. In economic forecasting [3], and environmental studies [4], precise measurements at a specific time period are unreliable, leading researchers to use the mean and confidence interval of measurements over time. Similarly, in engineering [5], multiple measurements of a single statistic are taken, and the mean along with the confidence interval is considered as the target. These uncertainties motivate the study of Interval Regression.
The difference between Interval Regression and Standard Regression
In the Standard Regression setting, each instance is associated with a single target value. In Interval Regression setting, each instance is associated with an interval of acceptable values, called target interval. There are four types of target intervals $(y_{l},y_{u})$: uncensored ($-\infty<y_{l}=y_{u}<\infty$), right-censored ($-\infty<y_{l}<y_{u}=\infty$), left-censored ($-\infty=y_{l}<y_{u}<\infty$), and interval-censored ($-\infty<y_{l}<y_{u}<\infty$). In this Interval Regression setting, for each instance, the model aims to predict a single value that falls within its target interval. Refer to the top two figures in Figure 1 for a clearer visual understanding. Standard Regression can be viewed as a special case of Interval Regression limited to uncensored target intervals.
Clarification
Although many studies are titled with “Interval Regression”, the setting of this study may differ from those works. For example, several studies such as [6, 7] focus on predicting an interval for each instance. Instead, in this study’s Interval Regression setting, the model predicts a single value for each instance, with the aim of falling within its target interval.
Literature Review
Several models have been specifically developed for this study’s Interval Regression setting, including Linear models [8, 9] and Tree models [10, 11]. Details on how these models operate can be found in Section 2.
Contribution
This study proposes models such as K-Nearest Neighbors (KNN), Random Forest, and deep neural networks, particularly multi-layer perceptron (MLP). It then offers a comprehensive comparison of the performance of Interval Regression models, including both existing models like Linear and Tree models, as well as the proposed ones, across a variety of datasets, including real-world and synthetic.
Figure 1: Example of converting Interval Regression into Standard Regression. In approach 1, each target interval is represented by two endpoints, while in approach 2, the interval is represented by the midpoint. The goal of Interval Regression is to predict a value falls within the target interval. So this example shows that these conversion approaches perform poorly in the setting of Interval Regression, that’s why they are not recommended.
2 Existing and Proposed Models
Notation
In the supervised Interval Regression setting, an instance is represented as $[\mathbf{x},(y_{l},y_{u})]$ where $\mathbf{x}\in\mathbf{R}^{m}$ is the feature vector and $(y_{l},y_{u})\in\mathbf{R}^{2}$, where $y_{l}\leq y_{u}$ (lower and upper target), is the target interval. The predicted value of a Interval Regression model is $\hat{y}\in\mathbf{R}$.
Noteworthy: Accelerated Failure Time (AFT) models
These models such as those in [12, 13, 14], are a family of models specifically designed for survival analysis. Survival analysis involves predicting the time until an event occurs, using interval-censored data where the exact event time is unknown but falls within a specified range. However, a limitation of these models is that they can only handle uncensored and right-censored intervals, which makes them unsuitable for the general setting of Interval Regression.
Noteworthy: Converting Interval Regression to Standard Regression
Instead of using the intervals as the targets, some studies represent the target interval by discrete points then use them as the Standard Regression targets. Drouin et al. [10] transform the interval into two points, $(\mathbf{x},y_{l})$ and $(\mathbf{x},y_{u})$, called Interval-CART model – Approach 1 in Figure 1, whereas Cheng et al. [9] convert it into its midpoint, $(\mathbf{x},\frac{y_{l}+y_{u}}{2})$ – Approach 2 in Figure 1. While thess approaches seem straightforward and reasonable, it overlooks the full potential of interval information by reducing intervals to finite values and fails to handle left-censored or right-censored intervals, resulting in wasted data. Consequently, the performance of these models is worse compared to Interval Regression models specifically designed for interval targets [10, 9].
2.1 Existing Models
Linear
The Linear model called Max Margin Interval Regression was introduced by Rigaill et al. [8]. The prediction formula is given by:
$\hat{y}=\mathbf{x}\cdot\beta+\beta_{0}$
where $\beta$ and $\beta_{0}$ are the parameters to be learned. To estimate these parameters, the authors extended the Standard Regression loss function—specifically, the Squared Error or Mean Squared Error (MSE)—into a more generalized form called the Hinge Squared Error. The error between prediction $\hat{y}$ and target interval $(y_{l},y_{u})$ is defined as follows:
$l\Big{(}\hat{y},(y_{l},y_{u})\Big{)}=\Big{(}\text{ReLU}(y_{l}-\hat{y}+\epsilon)\Big{)}^{p}+\Big{(}\text{ReLU}(\hat{y}-y_{u}+\epsilon)\Big{)}^{p}$ (1)
Here, $\epsilon>0$ is the margin length ($\epsilon=1$ by default) and the loss type $p=2$ are chosen. Cheng et al. [9] proposed a model called Regression with Interval Targets (RIT), which is essentially the same as the above model, but uses a loss type of $p=1$ instead.
Tree
The Tree model, called Maximum Margin Interval Trees (MMIT), was introduced by Drouin et al. [10] as a nonlinear model to Interval Regression. The prediction is given by:
$\hat{y}=\mathcal{T}(\mathbf{x})$
where $\mathcal{T}$ represents the tree model. The tree architecture follows the same structure as the regression decision tree CART (Classification And Regression Tree) [15]. The only difference lies in the regression value for each leaf: instead of taking the mean of all targets like CART does, it chooses a constant value that minimizes the mean Hinge Error (1) between this constant and the target intervals. They also introduce an alternative Interval Regression loss function, called Hinge Linear Error, which is defined by Equation (1) with $p=1$.
AFT Model in XGBoost
The AFT model in XGBoost, introduced by Barnwal et al. [11], is an ensemble tree-based model designed for survival analysis. The model makes predictions as follows:
$\hat{y}=\mathbf{T}(\mathbf{x}),$
where $\mathbf{T}$ represents the ensemble of trees in XGBoost. Although it belongs to the AFT model family, this model can handle all non-negative target intervals. To extend its applicability to real-valued target intervals $(y_{l},y_{u})\in\mathbf{R}^{2}$, Barnwal et al. [11] apply an exponential transformation to the original interval:
$(y_{l},y_{u})\longrightarrow(\exp y_{l},\exp y_{u}).$
After making a prediction $y_{p}$ in the exponential space, the result is mapped back to the original scale using the logarithm transformation:
$\hat{y}=\log y_{p}$
2.2 Proposed Models
K-Nearest Neighbors
KNN is a classical model for Standard Regression, which motivates its consideration in the Interval Regression setting. The Interval Regression procedure follows the same steps as the standard KNN regression model [16], with the key difference being how the regression value is determined from the $k$ nearest neighbors. We propose utilizing the MMIT regression function. Specifically, we treat the set of $k$ nearest neighbors as a single leaf in a MMIT, where the regression value is a constant that minimizes the mean Hinge Error (1) between this constant and the $k$ target intervals. The prediction $\hat{y}$ based on the target intervals of the $k$ nearest neighbors, denoted as $\{(y_{l}^{i},y_{u}^{i})\}_{i=1}^{k}$, is formulated as:
$\hat{y}=\operatorname*{argmin}_{c}\frac{1}{k}\sum_{i=1}^{k}l\big{(}c,(y_{l}^{i},y_{u}^{i})\big{)}$
Random Forest
The idea of using ensembles of MMITs is introduced in Chapter 6 (Discussion and Conclusions) of [10]. Since MMIT is essentially a decision tree regressor for the Interval Regression setting, its extension to construct a Random Forest for Interval Regression is straightforward. This proposed model is referred to as the Maximum Margin Interval Forest (MMIF), as it is composed of multiple MMITs. The operation of MMIF follows the same principles as the standard Random Forest regressor [17], with the key distinction being that MMITs are used instead of conventional decision trees.
MLP
Since MLP is a commonly used neural network and serves as a universal approximator in Standard Regression models [18], its application in Interval Regression is worth considering. MLP models for Interval Regression are proposed, where the Squared Hinge Error (1) is used for training.
3 Experiments
Experimental Datasets
To ensure a fair and comprehensive comparison across different scenarios, we evaluate all the models on a diverse set of datasets. Specifically, we use 36 real-world datasets from the UCI repository [19], along with 3 simulated datasets, identical to those in [10]. Each simulated dataset consists of 200 instances with 20 features, where 19 are noise and one is the true feature $x$. The target intervals are defined based on the linear function of $x$, $\sin(x)$, or the absolute value of $x$. All datasets are available at https://github.com/aldro61/mmit-data.
Train/Test Setup
Each dataset is divided into 5 similar sized folds. Each fold is used as the test set, while the remaining 4 folds are combined to form the train set.
Evaluation Metrics
For each train/test pair, the Interval Regression model is trained on the train set and used to make predictions on the test set. The test set is denoted as having $m$ instances with corresponding target intervals $\{(y_{l}^{i},y_{u}^{i})\}_{i=1}^{m}$, and the set of predictions is $\{\hat{y}^{i}\}_{i=1}^{m}$. The Mean Squared Hinge Error is given by:
$\frac{1}{m}\sum_{i=1}^{m}\left(\text{ReLU}(y_{l}^{i}-\hat{y}^{i})\right)^{2}+\left(\text{ReLU}(\hat{y}^{i}-y_{u}^{i})\right)^{2}$
For each dataset, 5 train/test pairs are used, resulting in 5 error values. From these 5 error values, the mean and standard deviation are computed. Two metrics are then considered: performance and consistency. A smaller mean error indicates better performance, while a smaller standard deviation indicates greater consistency.
3.1 Model Configuration
Except for the AFT model in XGBoost, the Squared Hinge Error is used as the evaluation function for both the cross-validation process and model training. All models are implemented using Python, except for the Linear model. The configuration for each model below is applied to a single train/test set pair.
Constant (Featureless)
This model predicts a single value from the train set using only target intervals. The value is a constant that minimizes the Squared Hinge Error (1) between this constant and the set of target intervals.
Linear
The R package penaltyLearning [20] is used to implement the Max Margin Interval Regression model with L1 regularization. The regularization parameter starts at 0.001, increasing by a factor of 1.2 until no features remain, with cross-validation (cv = 5) on the train set to determine the optimal L1 regularization value.
MMIT
[…]
AFT model in XGBoost
The model was implemented using the Python package xgboost. Cross-validation (cv = 5) was performed to select the optimal hyperparameters, following the same grid search described in Barnwal et al. [11]:
  • learning_rate: 0.001, 0.01, 0.1, 1.0
  • max_depth: 2, 3, 4, 5, 6, 7, 8, 9, 10
  • min_child_weight: 0.001, 0.1, 1.0, 10.0, 100.0
  • reg_alpha: 0.001, 0.01, 0.1, 1.0, 10.0, 100.0
  • reg_lambda: 0.001, 0.01, 0.1, 1.0, 10.0, 100.0
  • aft_loss_distribution_scale: 0.5, 0.8, 1.1, 1.4, 1.7, 2.0
KNN
The Euclidean distance metric is applied to normalized features (mean 0, standard deviation 1) to determine the nearest neighbors, implemented using the NearestNeighbors module from the sklearn.neighbors package. A cross-validation (cv = 5) is used to select the optimal value of $K$. The candidate values for $K$ range from 1 to $\lceil\sqrt{n}\rceil$, where $n$ is the number of train instances.
MMIF
MMIF is an ensemble of 100 MMITs. For each MMIT, two-thirds of the train dataset is randomly selected for training, while the remaining one-third serves as the out-of-bag (OOB) set for validation. Additionally, one-third of the original features are randomly chosen for each MMIT. A cross-validation (cv = 5) is used to select the optimal hyperparameters for each MMIT including max_depth {2, 5, 10, 15, 20, 25} and min_split_samples {2, 5, 10, 20, 50}. Each MMIT has its own OOB error, which is used to determine the aggregation rule. Let $\mathcal{T}_{i}$ be an MMIT with OOB error $e_{i}$. The weight assigned to $\mathcal{T}_{i}$ is given by $w_{i}=\frac{\frac{1}{e_{i}}}{\sum_{j=1}^{100}\frac{1}{e_{j}}}$ and the final prediction from MMIF is computed as $\sum_{i=1}^{100}w_{i}\mathcal{T}_{i}(\mathbf{x})$ where $\mathbf{x}$ is the set of features.
MLP
A cross-validation (cv = 5) is used to determine the optimal hyperparameters:
  • num_layer: {1, 2}
  • hidden_layer_size: {5, 10, 20}
  • activation: {ReLU, Sigmoid}
The model is trained using the Adam optimizer with a fixed learning rate of 0.001.
Figure 2: The mean and standard deviation of the log of test squared hinge errors from simulated datasets. The Linear model performs best when the dataset is linear. In nonlinear datasets, Tree-based models achieve the best performance.
Figure 3: The mean and standard deviation of test squared hinge errors for datasets with a high number of features (due to significantly higher error, the Constant model is omitted to improve comparative visualization). Tree-based models generally perform exceptionally well due to their inherent feature selection mechanism. On the other hand, while MLP with ReLU activation function is a more generalized Linear model, it fails to outperform the Linear model. One reason for this is that when a dataset contains a majority of noisy features, MLP cannot effectively reduce their impact on predictions in the same way that a Linear model with L1 regularization can.
Figure 4: The mean and standard deviation of test squared hinge errors for datasets with a moderate number of features. In these scenarios, MLP generally performs well.
Performance Consistency
1st 2nd 3rd 4th 5th 6th 7th 1st 2nd 3rd 4th 5th 6th 7th
constant 2 2 3 1 4 7 20 4 0 2 3 3 8 19
knn 4 4 2 6 6 16 1 4 3 3 7 7 12 3
linear 3 11 6 8 5 5 1 7 7 8 7 7 3 0
mmit 2 10 9 3 9 4 2 2 13 4 8 8 2 2
mmif 15 5 8 7 4 0 0 12 9 9 4 3 2 0
aft_xgb 4 4 3 4 7 5 12 4 1 7 2 3 10 12
mlp 9 3 8 10 4 2 3 6 6 6 8 8 2 3
Table 1: The comparison of the performance and consistency of 7 models across 39 datasets.
4 Discussion
Dataset Variations
To ensure a fair and comprehensive comparison, the quality of the datasets is crucial. First, the quality of features is considered. The Constant model (or Featureless), being the simplest, is expected to perform the worst across all datasets. However, in 19 datasets, it is not ranked the lowest. Notably, in 8 of these datasets, it is ranked 4th place or higher, suggesting that the features in these datasets provide limited predictive value.
Next, the relationship between features and targets is examined. Since Linear models assume a linear relationship, while MMIT assumes a highly nonlinear one, dataset linearity is assessed based on performance and consistency from both Linear and MMIT models. If the Linear model outperforms MMIT with greater consistency, the dataset is more likely to exhibit a linear relationship. However, datasets where Linear performs better but with lower consistency than MMIT are not considered. From that criteria, out of 39 datasets, 15 are identified as more linear than nonlinear, including autohorse, autompg, bodyfat, breasttumor, cholesterol, cleveland, elusage, fishcatch, fruity, meta, pbc, simulated.linear, sleep, and vineyard. Conversely, 15 datasets are more nonlinear than linear, such as autoprice, basketball, cloud, cpu, lowbwt, lymphoma.mkatayama, machine.cpu, mbagrade, pollution, pwlinear, pyrim, servo, simulated.abs, simulated.sin, and strike. This statistic demonstrates the diversity and quality of the datasets, ensuring a fair comparison across different scenarios.
Comprehensive Comparison
Let’s examine the performance and consistency rankings in Table 1. Starting with performance, it’s clear that MMIF outperforms other models in most cases. Linear, MMIT and MLP are comparable, while Constant, KNN, and the AFT model in XGBoost show the poorest performance. When it comes to consistency, MMIF stands out as the most consistent. This is expected, as MMIF being an ensemble model, relies on a group of weak MMITs to make the final prediction, which helps improve its stability. Again, Constant, KNN, and the AFT model in XGBoost show the lowest consistency.
No single model is optimal for all scenarios
As shown in Table 1, there is no universally optimal model for every scenario. However, if complexity is not a primary concern, and the goal is to select the model with the best overall performance and consistency, MMIF is likely the first choice. Alternatively, if model complexity and training time are critical factors, MMIT or Linear are preferable, as they are less complex than MMIF, require significantly fewer training resources, and still maintain reliable performance.
MMIT vs MMIF
This is a comparison worth considering. MMIF was introduced with the expectation that an ensemble of MMITs would outperform a single MMIT, and the results confirm this hypothesis. Specifically, across 39 datasets, MMIF is found to outperform MMIT in 28 of them, which constitutes the majority. The reason for this is straightforward: MMIF, being an ensemble of MMITs, is able to reduce overfitting by not relying on a single MMIT.
Discussion on the AFT Model in XGBoost and potential future work
The AFT model in XGBoost is one of the most complex models in this study, but its performance and consistency do not justify the complexity. It has a limitation in that it can only handle non-negative target intervals. To make the model applicable to a general Interval Regression setting, the target intervals must be transformed using exponential function. This introduces a weakness, as the model becomes highly sensitive when the dataset contains many left-censored target intervals. When predictions are made on the exponential scale, if the predicted value is close to zero, the mapping back to the original scale can result in large negative values, as seen in some test cases where predictions like $-10^{6}$ were made.
[…]
The experiments conducted in [11] indicate that the AFT model in XGBoost performs among the worst across their six simulated datasets when compared to the Linear model and MMIT.
A potential direction for future work is to transform left-censored intervals into interval-censored intervals by replacing all lower bounds of $\infty$- ∞ with a finite value. This adjustment aims to reduce the model’s sensitivity to left-censored intervals and mitigate the tendency to produce excessively large negative predictions.
Code Available
All experiment code is available in this GitHub repository: https://github.com/lamtung16/ML_IntervalRegression for reproducibility and further exploration of the models used in this study.
SUPPLEMENTAL MATERIALS
[…]
constant knn linear mmit mmif aft_xgb mlp
auto93 7 6 3 5 1 2 4
autohorse 7 4 2 5 1 6 3
autompg 7 6 2 3 4 5 1
autoprice 7 4 6 2 1 5 3
baskball 6 2 5 3 1 7 4
bodyfat 7 6 2 4 3 5 1
breasttumor 6 1 4 5 2 7 3
cholesterol 3 5 4 6 1 7 2
cleveland 6 4 2 5 3 7 1
cloud 7 6 3 2 5 4 1
cpu 7 6 4 3 1 2 5
echomonths 7 4 2 5 3 6 1
elusage 7 2 1 3 4 5 6
fishcatch 7 6 4 5 3 2 1
fruitfly 1 5 2 6 3 7 4
housing 7 3 6 5 4 2 1
lowbwt 6 5 3 2 1 7 4
lymphoma.mkatayama 1 5 6 3 4 7 2
lymphoma.tdh 3 6 2 4 1 7 5
machine.cpu 7 6 5 3 2 1 4
mbagrade 3 6 5 1 2 4 7
meta 5 1 2 6 4 7 3
pbc 7 5 1 4 2 6 3
pharynx 7 6 2 1 3 4 5
pollution 4 6 3 2 1 7 5
pwlinear 7 6 5 2 3 4 1
pyrim 2 1 6 3 4 5 7
sensory 2 6 3 5 1 7 4
servo 7 6 4 2 5 1 3
simulated.abs 5 4 6 2 1 3 7
simulated.linear 7 6 1 3 2 5 4
simulated.sin 5 7 4 2 1 3 6
sleep 7 1 2 5 3 6 4
stock 7 2 5 3 1 6 4
strike 6 2 4 7 5 1 3
triazines 5 6 7 2 1 3 4
veteran 6 3 2 7 4 5 1
vineyard 7 4 3 6 5 1 2
wisconsin 6 5 4 2 1 7 3
Table 2: The performance ranking of 7 models across 39 datasets.
constant knn linear mmit mmif aft_xgb mlp
auto93 7 6 4 2 1 3 5
autohorse 7 4 2 5 1 6 3
autompg 7 5 1 4 2 6 3
autoprice 7 3 4 2 1 6 5
baskball 5 4 3 2 1 6 7
bodyfat 7 5 1 4 3 6 2
breasttumor 6 1 2 5 3 7 4
cholesterol 5 6 4 7 1 3 2
cleveland 6 5 1 4 3 7 2
cloud 7 6 3 2 5 4 1
cpu 7 4 5 2 1 3 6
echomonths 7 2 5 4 3 6 1
elusage 7 4 1 2 3 5 6
fishcatch 7 6 2 5 4 3 1
fruitfly 3 5 2 7 1 6 4
housing 7 3 4 5 2 6 1
lowbwt 6 4 3 1 2 7 5
lymphoma.mkatayama 1 6 5 2 3 7 4
lymphoma.tdh 3 6 4 2 1 7 5
machine.cpu 7 6 5 2 3 1 4
mbagrade 6 2 5 1 3 4 7
meta 6 1 2 5 4 7 3
pbc 6 2 3 5 4 7 1
pharynx 6 5 1 3 4 7 2
pollution 4 6 3 2 1 7 5
pwlinear 7 6 5 4 2 3 1
pyrim 1 3 6 4 5 2 7
sensory 1 6 3 4 2 7 5
servo 7 5 4 2 6 1 3
simulated.abs 6 7 5 2 1 3 4
simulated.linear 7 6 1 4 2 5 3
simulated.sin 4 7 6 2 1 3 5
sleep 7 4 2 5 1 6 3
stock 7 1 2 5 3 6 4
strike 5 7 6 3 2 1 4
triazines 1 5 3 6 2 7 4
veteran 7 1 4 3 6 5 2
vineyard 7 4 3 6 5 1 2
wisconsin 4 6 1 3 2 7 5
Table 3: The consistency ranking of 7 models across 39 datasets.

### v2

@@
 Abstract
-Regression models are essential for a wide range of real-world applications.
-However, in practice, target values are not always precisely known; instead, they may be represented as intervals of acceptable values.
+Regression models are common for a wide range of real-world applications.
+However, for the data collection in practice, target values are not always precisely known; instead, they may be represented as intervals of acceptable values.
 This challenge has led to the development of Interval Regression models.
-In this study, we provide a comprehensive review of existing Interval Regression models and introduce alternative models for comparative analysis.
+In this study, we provide a comprehensive review of existing Interval Regression models and propose alternative models for comparative analysis.
 Experiments are conducted on both real-world and synthetic datasets to offer a broad perspective on model performance.
@@
 keywords:
-interval regression, survival analysis, supervised machine learning.
+regression, interval data, survival analysis, supervised machine learning.
 1 Introduction
-Need for Interval Regression over Standard Regression
 Regression models are widely used in real-world applications.
-However, due to some reasons, such as errors in measurement, the data collection procedures, or human factors, the target value in many scenarios is not a single value but rather an interval of acceptable values.
-For instance, in medical research [1, 2], the precise survival time of a patient is often uncertain, but it is known that the patient survives for at least a certain number of years.
-In economic forecasting [3], and environmental studies [4], precise measurements at a specific time period are unreliable, leading researchers to use the mean and confidence interval of measurements over time.
-Similarly, in engineering [5], multiple measurements of a single statistic are taken, and the mean along with the confidence interval is considered as the target.
+However, due to some reasons (errors in measurement, data collection procedures or human factors), the target value is not a single value but rather an interval of acceptable values.
+One example is in survival analysis [1, 2], the exact survival time of a patient is often uncertain, though it is typically known that the patient has survived for at least a certain number of years.
+In economic forecasting [3], and environmental studies [4], precise measurements at a specific time period are unreliable, leading researchers to use the confidence interval of measurements over time.
+Similarly, in engineering [5], multiple measurements of a single statistic are taken, and the confidence interval is considered as the target.
 These uncertainties motivate the study of Interval Regression.
-The difference between Interval Regression and Standard Regression
-In the Standard Regression setting, each instance is associated with a single target value.
-In Interval Regression setting, each instance is associated with an interval of acceptable values, called target interval.
-There are four types of target intervals $(y_{l},y_{u})$: uncensored ($-\infty<y_{l}=y_{u}<\infty$), right-censored ($-\infty<y_{l}<y_{u}=\infty$), left-censored ($-\infty=y_{l}<y_{u}<\infty$), and interval-censored ($-\infty<y_{l}<y_{u}<\infty$).
-In this Interval Regression setting, for each instance, the model aims to predict a single value that falls within its target interval.
-Refer to the top two figures in Figure 1 for a clearer visual understanding.
-Standard Regression can be viewed as a special case of Interval Regression limited to uncensored target intervals.
-Clarification
-Although many studies are titled with “Interval Regression”, the setting of this study may differ from those works.
-For example, several studies such as [6, 7] focus on predicting an interval for each instance.
-Instead, in this study’s Interval Regression setting, the model predicts a single value for each instance, with the aim of falling within its target interval.
-Literature Review
-Several models have been specifically developed for this study’s Interval Regression setting, including Linear models [8, 9] and Tree models [10, 11].
-Details on how these models operate can be found in Section 2.
-Contribution
-This study proposes models such as K-Nearest Neighbors (KNN), Random Forest, and deep neural networks, particularly multi-layer perceptron (MLP).
-It then offers a comprehensive comparison of the performance of Interval Regression models, including both existing models like Linear and Tree models, as well as the proposed ones, across a variety of datasets, including real-world and synthetic.
+training dataset
+features (predictors) target (ground truth) prediction (response)
+Standard Regression $\mathbf{x}\in\mathbf{R}^{m}$ $y\in\mathbf{R}$ $\hat{y}\in\mathbf{R}$
+Interval Regression $\mathbf{x}\in\mathbf{R}^{m}$ $(y_{l},y_{u})\in\mathbf{R}^{2},y_{l}\leq y_{u}$ $\hat{y}\in\mathbf{R}$
+Table 1: Comparison of Standard Regression versus Interval Regression setting.
+The term “regression” typically refers to Standard Regression (or Point Regression), where each instance is associated with a single target value.
+This study focuses on Interval Regression setting, where each instance is associated with an interval of acceptable values, called target interval.
+There are 4 types of target intervals $(y_{l},y_{u})$ based on the end values: uncensored ($\infty<y_{l}=y_{u}<\infty$), right-censored ($\infty<y_{l}<y_{u}=\infty$), left-censored ($\infty=y_{l}<y_{u}<\infty$), and interval-censored ($\infty<y_{l}<y_{u}<\infty$).
+Although many studies are titled with “Interval Regression”, such as [6, 7, 8, 9], the setting of these studies differs from this work since they focus on predicting an interval for each instance.
+Instead, in this study’s Interval Regression setting, the model predicts a single value for each instance, see Table 1.
+In some studies, Interval Regression models are trained by converting the problem into Standard Regression: instead of using intervals as targets, the target interval is represented by discrete points, which are then treated as Standard Regression targets.
+Using the same notations in Table 1 for $\mathbf{x}$ as the features set and $(y_{l},y_{u})$ as the target interval, Drouin et al.
+[10] transform the instance $\big(\mathbf{x},(y_{l},y_{u})\big)$ into two points, $(\mathbf{x},y_{l})$ and $(\mathbf{x},y_{u})$, called Interval-CART model whereas Cheng et al.
+[11] convert it into its midpoint, $(\mathbf{x},\frac{y_{l}+y_{u}}{2})$ – Approach 1 and 2 in Figure 1.
+While these approaches are reasonable, it overlooks the interval information by representing intervals by finite values and fails to handle left-censored or right-censored intervals, resulting in wasted data.
+So the performance of these models is worse compared to the models specifically designed for interval targets, see the results in [10, 11] and Figure 1 example.
+To date, existing Interval Regression models for this setting include Linear [12, 11, 13], Tree-based [10, 14] and Accelerated Failure Time (AFT) models family [15, 16, 17, 18].
+Details on these models can be found in Section 2.
+Contribution.
+This study proposes some alternative Interval Regression models and offers a comprehensive performance comparison of all models (existing and proposed) across a variety of datasets, including real-world and synthetic.
 Figure 1: Example of converting Interval Regression into Standard Regression.
-In approach 1, each target interval is represented by two endpoints, while in approach 2, the interval is represented by the midpoint.
+In approach 1, each interval instance is represented by two endpoints, while in approach 2, it is represented by the midpoint.
 The goal of Interval Regression is to predict a value falls within the target interval.
-So this example shows that these conversion approaches perform poorly in the setting of Interval Regression, that’s why they are not recommended.
-2 Existing and Proposed Models
-Notation
-In the supervised Interval Regression setting, an instance is represented as $[\mathbf{x},(y_{l},y_{u})]$ where $\mathbf{x}\in\mathbf{R}^{m}$ is the feature vector and $(y_{l},y_{u})\in\mathbf{R}^{2}$, where $y_{l}\leq y_{u}$ (lower and upper target), is the target interval.
-The predicted value of a Interval Regression model is $\hat{y}\in\mathbf{R}$.
-Noteworthy: Accelerated Failure Time (AFT) models
-These models such as those in [12, 13, 14], are a family of models specifically designed for survival analysis.
-Survival analysis involves predicting the time until an event occurs, using interval-censored data where the exact event time is unknown but falls within a specified range.
-However, a limitation of these models is that they can only handle uncensored and right-censored intervals, which makes them unsuitable for the general setting of Interval Regression.
-Noteworthy: Converting Interval Regression to Standard Regression
-Instead of using the intervals as the targets, some studies represent the target interval by discrete points then use them as the Standard Regression targets.
-Drouin et al.
-[10] transform the interval into two points, $(\mathbf{x},y_{l})$ and $(\mathbf{x},y_{u})$, called Interval-CART model – Approach 1 in Figure 1, whereas Cheng et al.
-[9] convert it into its midpoint, $(\mathbf{x},\frac{y_{l}+y_{u}}{2})$ – Approach 2 in Figure 1.
-While thess approaches seem straightforward and reasonable, it overlooks the full potential of interval information by reducing intervals to finite values and fails to handle left-censored or right-censored intervals, resulting in wasted data.
-Consequently, the performance of these models is worse compared to Interval Regression models specifically designed for interval targets [10, 9].
+This example shows that these conversion approaches perform poorly in Interval Regression setting, so they are not recommended.
+2 Interval Regression Models
+This section details Interval Regression models, including existing and proposed ones.
+See Table 2 for all of the employed models in this study.
+The notation used hereafter follows Table 1, where $\mathbf{x}\in\mathbf{R}^{m}$ is the model input (a vector of $m$ features), $(y_{l},y_{u})\in\mathbf{R}^{2}$ where $y_{l}\leq y_{u}$ is the target interval, and $\hat{y}\in\mathbf{R}$ is the model prediction.
 2.1 Existing models
-Linear
-The Linear model called Max Margin Interval Regression was introduced by Rigaill et al. [8]. The prediction formula is given by:
-$\hat{y}=\mathbf{x}\cdot\beta+\beta_{0}$
-where $\beta$ and $\beta_{0}$ are the parameters to be learned.
-To estimate these parameters, the authors extended the Standard Regression loss function—specifically, the Squared Error or Mean Squared Error (MSE)—into a more generalized form called the Hinge Squared Error.
-The error between prediction $\hat{y}$ and target interval $(y_{l},y_{u})$ is defined as follows:
+Max Margin Interval Regression (linear)
+This linear model was introduced by Rigaill et al. [12], and its extended theoretical foundations were later studied by Cheng et al. [11]. The prediction formula is given by:
+$\hat{y}=\mathbf{x}\cdot\beta+\beta_{0}$ (1)
+where $\beta\in\mathbf{R}^{m}$ and $\beta_{0}\in\mathbf{R}$ are the parameters.
+To estimate these parameters, the authors extended the Standard Regression loss function — the Squared Error (SE) — into a generalized form called the Hinge Squared Error.
+The error value between prediction $\hat{y}$ and target interval $(y_{l},y_{u})$ is defined as:
+model name source code programming language citation
+linear Hocking [19] R Rigaill et al. [12]
+mmit Drouin [20] Python binding C++ Drouin et al. [10]
+aft_xgb Chen and Guestrin [21] Python Barnwal et al. [14]
+constant paper GitHub repo Python binding C++ additional baseline
+knn paper GitHub repo Python proposed
+mlp paper GitHub repo Python proposed
+mmif paper GitHub repo Python binding C++ proposed
+Table 2: List of the employed models in this study with corresponding source.
+$l\Big(\hat{y},(y_{l},y_{u})\Big)=\begin{cases}(y_{l}-\hat{y}+\epsilon)^{2},&\mbox{if }\hat{y}<y_{l}+\epsilon\\ 0,&\mbox{if }y_{l}+\epsilon\leq\hat{y}\leq y_{u}-\epsilon\\ (\hat{y}-y_{u}+\epsilon)^{2},&\mbox{if }\hat{y}>y_{u}-\epsilon\end{cases}$
+where $\epsilon\geq 0$ is the margin length. This error becomes SE when $y_{l}+\epsilon=y_{u}-\epsilon$.
+Cheng et al. [11] proposed a model called Regression with Interval Targets (RIT), which is essentially the same as model (1), but uses Hinge Absolute Error instead:
+$l\Big(\hat{y},(y_{l},y_{u})\Big)=\begin{cases}|y_{l}-\hat{y}+\epsilon|,&\mbox{if }\hat{y}<y_{l}+\epsilon\\ 0,&\mbox{if }y_{l}+\epsilon\leq\hat{y}\leq y_{u}-\epsilon\\ |\hat{y}-y_{u}+\epsilon|,&\mbox{if }\hat{y}>y_{u}-\epsilon\end{cases}$
+The general Interval Regression loss function can be written in this short form:
 $l\Big(\hat{y},(y_{l},y_{u})\Big)=\Big(\text{ReLU}(y_{l}-\hat{y}+\epsilon)\Big)^{p}+\Big(\text{ReLU}(\hat{y}-y_{u}+\epsilon)\Big)^{p}$ (2)
-Here, $\epsilon>0$ is the margin length ($\epsilon=1$ by default) and the loss type $p=2$ are chosen.
-Cheng et al.
-[9] proposed a model called Regression with Interval Targets (RIT), which is essentially the same as the above model, but uses a loss type of $p=1$ instead.
-Tree
-The Tree model, called Maximum Margin Interval Trees (MMIT), was introduced by Drouin et al. [10] as a nonlinear model to Interval Regression. The prediction is given by:
+where $p\in\{1,2\}$. See Figure 2 for the visual comparison between hinge errors and SE.
+To estimate the parameters $\beta$ and $\beta_{0}$ in (1), apply a gradient descent algorithm with respect to $\beta$ and $\beta_{0}$ to minimize the sum of Hinge Error (2) over the training set.
+Rigaill et al.
+[12] employed the Fast Iterative Shrinkage-Thresholding Algorithm (FISTA) [22], while Cheng et al.
+[11] employed stochastic gradient descent (SGD).
+Maximum Margin Interval Trees (MMIT)
+This Tree model was introduced by Drouin et al. [10] as an alternative nonlinear model to Max Margin Interval Regression. The prediction is given by:
 $\hat{y}=\mathcal{T}(\mathbf{x})$
 where $\mathcal{T}$ represents the tree model.
-The tree architecture follows the same structure as the regression decision tree CART (Classification And Regression Tree) [15].
-The only difference lies in the regression value for each leaf: instead of taking the mean of all targets like CART does, it chooses a constant value that minimizes the mean Hinge Error (1) between this constant and the target intervals.
-They also introduce an alternative Interval Regression loss function, called Hinge Linear Error, which is defined by Equation (1) with $p=1$.
-AFT Model in XGBoost
+The tree has the same structure as CART (Classification And Regression Tree) [23].
+The only difference lies in the regression value for each leaf which is a set of $S$ instances $\{\mathbf{x}^{i},(y_{l}^{i},y_{u}^{i})\}_{i=1}^{S}$: instead of taking the mean of all targets like CART does, it chooses a constant value $c$ that minimizes the mean Hinge Error (2) between this constant and the target intervals:
+$\hat{y}=\operatorname*{argmin}_{c}\frac{1}{S}\sum_{i=1}^{S}l\big(c,(y_{l}^{i},y_{u}^{i})\big),\quad c\in\mathbf{R}$
+A heuristic algorithm is used to speed up the estimation of the mean value for each leaf by considering a finite set of candidates, rather than solving for the optimal $c$:
+$\hat{y}=\operatorname*{argmin}_{c}\frac{1}{S}\sum_{i=1}^{S}l\big(c,(y_{l}^{i},y_{u}^{i})\big),\quad c\in\{y_{l}^{i},y_{u}^{i},\frac{y_{l}^{i}+y_{u}^{i}}{2}|y_{l}^{i}>-\infty,y_{u}^{i}<\infty\}_{i=1}^{S}$ (3)
+Figure 2: Visualization of the Loss Functions: Error values relative to predictions and targets.
+The training procedure follows the same operations as in CART to minimize the sum of Hinge Error (2) over the training set.
+AFT Model in XGBoost (aft_xgb)
 The AFT model in XGBoost, introduced by Barnwal et al. [14], is an ensemble tree-based model designed for survival analysis. The model makes predictions as follows:
-$\hat{y}=\mathbf{T}(\mathbf{x}),$
+$y_{p}=\mathbf{T}(\mathbf{x}),$
 where $\mathbf{T}$ represents the ensemble of trees in XGBoost.
-Although it belongs to the AFT model family, this model can handle all non-negative target intervals.
+This model can handle only non-negative target intervals.
 To extend its applicability to real-valued target intervals $(y_{l},y_{u})\in\mathbf{R}^{2}$, Barnwal et al.
-[11] apply an exponential transformation to the original interval:
-$(y_{l},y_{u})\longrightarrow(\exp y_{l},\exp y_{u}).$
+[14] apply an exponential transformation to the original interval.
 After making a prediction $y_{p}$ in the exponential space, the result is mapped back to the original scale using the logarithm transformation:
-$\hat{y}=\log y_{p}$
+$\hat{y}=\log y_{p}=\log\mathbf{T}(\mathbf{x})$
+Training this model is based on minimizing the negative log-likelihood function (loss function), which incorporates all training instances:
+$l_{AFT}\Big(\hat{y},(y_{l},y_{u})\Big)=\begin{cases}-\log\big(f_{Z}(\frac{\log y_{u}-\hat{y}}{\sigma})\times\frac{1}{\sigma y_{u}}\big)&\text{if $y_{l}=y_{u}$}\\ -\log\Big[F_{Z}\big(\frac{\log y_{u}-\hat{y}}{\sigma}\big)-F_{Z}\big(\frac{\log y_{l}-\hat{y}}{\sigma}\big)\Big]&\text{if $y_{l}<y_{u}$}\end{cases}$ (4)
+where $\sigma>0$ is the scale parameter. The Probability Density Function $f_{Z}(z)$ and Cumulative Distribution Function $F_{Z}(z)$ depends on the chosen distribution:
+  • normal: $f_{Z}(z)=\frac{\exp(-z^{2}/2)}{\sqrt{2\pi}}$, 00 $F_{Z}(z)=\frac{1}{2}\big(1+\text{erf}(\frac{z}{\sqrt{2}})\big)$
+  • logistic: $f_{Z}(z)=\frac{e^{z}}{(1+e^{z})^{2}}$, 0000 $F_{Z}(z)=\frac{e^{z}}{1+e^{z}}$
+  • extreme: $f_{Z}(z)=e^{z}e^{-\exp z}$, 0 $F_{Z}(z)=1-e^{-\exp z}$
+Figure 3 visualizes the loss function for each distributions, which can be compared to Figure 2 to observe their similarities.
+The model is trained using the gradient boosting algorithm [24] to minimize the loss value (4) over the training dataset given the distribution and the scale parameter $\sigma$.
+Figure 3: Visualization of the AFT XGBoost Loss (negative log-likelihood) for 3 distributions.
 2.2 Proposed Models
-K-Nearest Neighbors
-KNN is a classical model for Standard Regression, which motivates its consideration in the Interval Regression setting.
-The Interval Regression procedure follows the same steps as the standard KNN regression model [16], with the key difference being how the regression value is determined from the $k$ nearest neighbors.
+In addition to existing models, this study proposes several alternative models, including machine learning architectures such as KNN, Neural Networks and Forests, to provide a more comprehensive comparison.
+K-Nearest Neighbors (KNN)
+KNN is a classical model for Standard Regression, which motivates its consideration for Interval Regression.
+The operation follows the same steps as the standard KNN regression model [25], with the only difference being how the regression value is determined from the $k$ nearest neighbors.
 We propose utilizing the MMIT regression function.
@@
 The prediction $\hat{y}$ based on the target intervals of the $k$ nearest neighbors, denoted as $\{(y_{l}^{i},y_{u}^{i})\}_{i=1}^{k}$, is formulated as:
-$\hat{y}=\operatorname*{argmin}_{c}\frac{1}{k}\sum_{i=1}^{k}l\big{(}c,(y_{l}^{i},y_{u}^{i})\big{)}$
-Random Forest
-The idea of using ensembles of MMITs is introduced in Chapter 6 (Discussion and Conclusions) of [10].
-Since MMIT is essentially a decision tree regressor for the Interval Regression setting, its extension to construct a Random Forest for Interval Regression is straightforward.
-This proposed model is referred to as the Maximum Margin Interval Forest (MMIF), as it is composed of multiple MMITs.
-The operation of MMIF follows the same principles as the standard Random Forest regressor [17], with the key distinction being that MMITs are used instead of conventional decision trees.
-MLP
-Since MLP is a commonly used neural network and serves as a universal approximator in Standard Regression models [18], its application in Interval Regression is worth considering.
-MLP models for Interval Regression are proposed, where the Squared Hinge Error (1) is used for training.
+$\hat{y}=\operatorname*{argmin}_{c}\frac{1}{k}\sum_{i=1}^{k}l\big(c,(y_{l}^{i},y_{u}^{i})\big),\quad c\in\{y_{l}^{i},y_{u}^{i},\frac{y_{l}^{i}+y_{u}^{i}}{2}|y_{l}^{i}>-\infty,y_{u}^{i}<\infty\}_{i=1}^{k}$
+Multilayer Perceptron (MLP)
+Since MLP is a commonly used neural network in Standard Regression models [26], its application in Interval Regression is worth considering.
+The training process of this model is the same as in the linear model Max Margin Interval Regression.
+Maximum Margin Interval Forest (MMIF)
+The idea of using ensembles of MMITs was mentioned in Chapter 6 (Discussion and Conclusions) of [10].
+Since MMIT is a tree for the Interval Regression setting, its extension to construct a Forest is straightforward.
+The operation of MMIF follows the same principles as the standard Random Forest regressor [27], with the key distinction being that MMITs are used instead of CART.
+dataset features instances features pattern/quality
+linear nonlinear noisy excessive
+auto93 62 82 $\checkmark$ $\checkmark$
+autohorse 59 159 $\checkmark$ $\checkmark$
+autompg 25 392 $\checkmark$
+autoprice 15 159 $\checkmark$
+baskball 4 96 $\checkmark$
+bodyfat 14 252 $\checkmark$
+breasttumor 40 286 $\checkmark$
+cholesterol 26 299 $\checkmark$ $\checkmark$
+cleveland 26 299 $\checkmark$
+cloud 10 108 $\checkmark$
+cpu 36 209 $\checkmark$
+echomonths 13 106 $\checkmark$
+elusage 13 55 $\checkmark$
+fishcatch 15 158 $\checkmark$
+fruitfly 8 125 $\checkmark$ $\checkmark$
+housing 14 506 $\checkmark$
+lowbwt 23 189 $\checkmark$
+machine.cpu 6 209 $\checkmark$
+mbagrade 3 61 $\checkmark$ $\checkmark$
+meta 54 264 $\checkmark$ $\checkmark$
+pbc 29 276 $\checkmark$
+pollution 15 60 $\checkmark$
+pwlinear 10 200 $\checkmark$
+pyrim 27 74 $\checkmark$ $\checkmark$
+sensory 36 576 $\checkmark$ $\checkmark$
+servo 19 167 $\checkmark$
+sleep 7 51 $\checkmark$
+stock 9 950 $\checkmark$
+strike 23 625 $\checkmark$
+triazines 60 186 $\checkmark$ $\checkmark$
+veteran 13 137 $\checkmark$
+vineyard 3 52 $\checkmark$
+wisconsin 32 194 $\checkmark$
+simulated.abs 20 200 $\checkmark$ $\checkmark$
+simulated.linear 20 200 $\checkmark$ $\checkmark$
+simulated.sin 20 200 $\checkmark$ $\checkmark$
+pharynx 218 195 $\checkmark$ $\checkmark$
+lymphoma.mkatayama 258 41 $\checkmark$ $\checkmark$ $\checkmark$
+lymphoma.tdh 258 23 $\checkmark$ $\checkmark$ $\checkmark$
+Table 3: All datasets in this study.
+Linear means the pattern between features and targets is likely linear, while nonlinear indicates otherwise.
+Noisy refers to cases where more than half of the features are irrelevant, excessive describes the dataset having more features than instances.
+Figure 4: The mean and standard deviation of the log of test squared hinge errors from simulated datasets.
+The Linear model performs best when the dataset is linear.
+In nonlinear datasets, Tree-based models achieve the best performance.
 3 Experiments
 Experimental Datasets
-To ensure a fair and comprehensive comparison across different scenarios, we evaluate all the models on a diverse set of datasets.
-Specifically, we use 36 real-world datasets from the UCI repository [19], along with 3 simulated datasets, identical to those in [10].
+To ensure a comprehensive comparison across different scenarios, we evaluate all the models on a diverse set of datasets.
+Specifically, we use 36 real-world datasets from the UCI repository [28] including 3 biology datasets having excessive features, along with 3 simulated datasets - identical to those in [10].
 Each simulated dataset consists of 200 instances with 20 features, where 19 are noise and one is the true feature $x$.
 The target intervals are defined based on the linear function of $x$, $\sin(x)$, or the absolute value of $x$.
-All datasets are available at https://github.com/aldro61/mmit-data.
-Train/Test Setup
-Each dataset is divided into 5 similar sized folds. Each fold is used as the test set, while the remaining 4 folds are combined to form the train set.
+The real-valued outputs in these data sets were transformed into censored intervals, a small vertical shift was added by sampling a value from $\mathcal{N}(0,\frac{y_{l}}{10})$ and adding it to both interval bounds.
+All datasets are described in Table 3 and are accessible here: study datasets link.
 Evaluation Metrics
-For each train/test pair, the Interval Regression model is trained on the train set and used to make predictions on the test set.
-The test set is denoted as having $m$ instances with corresponding target intervals $\{(y_{l}^{i},y_{u}^{i})\}_{i=1}^{m}$, and the set of predictions is $\{\hat{y}^{i}\}_{i=1}^{m}$.
+Each dataset is divided into 5 similar sized folds.
+Each fold is used as the test set, while the remaining 4 folds are combined to form the train set.
+For each train/test pair, the model is trained on the train set and used to make predictions on the test set.
+The test set is denoted as having $M$ instances with corresponding target intervals $\{(y_{l}^{i},y_{u}^{i})\}_{i=1}^{M}$, and the predictions set is $\{\hat{y}^{i}\}_{i=1}^{M}$.
 The Mean Squared Hinge Error is given by:
-$\frac{1}{m}\sum_{i=1}^{m}\left(\text{ReLU}(y_{l}^{i}-\hat{y}^{i})\right)^{2}+\left(\text{ReLU}(\hat{y}^{i}-y_{u}^{i})\right)^{2}$
-For each dataset, 5 train/test pairs are used, resulting in 5 error values.
+$\frac{1}{M}\sum_{i=1}^{M}\left(\text{ReLU}(y_{l}^{i}-\hat{y}^{i})\right)^{2}+\left(\text{ReLU}(\hat{y}^{i}-y_{u}^{i})\right)^{2}$ (5)
+which is equivalent to using Hinge Error (2) with $p=2$ and $\epsilon=0$.
+Each train/test pair produces one error values.
 From these 5 error values, the mean and standard deviation are computed.
 Two metrics are then considered: performance and consistency.
-A smaller mean error indicates better performance, while a smaller standard deviation indicates greater consistency.
+A smaller mean error indicates better performance, while a smaller standard deviation indicates better consistency.
+Figure 5: The mean and standard deviation of test squared hinge errors for datasets with a high number of features.
+Tree-based models generally perform well due to their inherent feature selection mechanism.
+On the other hand, while MLP with ReLU activation function is a more generalized Linear model, it fails to outperform the Linear model.
+One reason for this is that when a dataset contains a majority of noisy features, MLP cannot effectively reduce their impact on predictions in the same way that a Linear model with L1 regularization can.
 3.1 Model Configuration
-Except for the AFT model in XGBoost, the Squared Hinge Error is used as the evaluation function for both the cross-validation process and model training.
-All models are implemented using Python, except for the Linear model.
+Except for the AFT model in XGBoost, the Squared Hinge Error is used as the loss function for both the cross-validation process and model training.
 The configuration for each model below is applied to a single train/test set pair.
@@
 This model predicts a single value from the train set using only target intervals.
-The value is a constant that minimizes the Squared Hinge Error (1) between this constant and the set of target intervals.
+The value is a constant that minimizes the Mean Squared Hinge Error (5) between this constant and the set of target intervals.
 Linear
-The R package penaltyLearning [20] is used to implement the Max Margin Interval Regression model with L1 regularization.
-The regularization parameter starts at 0.001, increasing by a factor of 1.2 until no features remain, with cross-validation (cv = 5) on the train set to determine the optimal L1 regularization value.
+The model Max Margin Interval Regression model is implemented with L1 regularization.
+The regularization parameter starts at 0.001, geometrically increasing by a factor of 1.2 until no features remain, with cross-validation (cv = 5) on the train set to determine the optimal L1 value.
 MMIT
-[…]
+A cross-validation (cv = 5) was used to select the optimal hyperparameters, including max_depth {0, 1, 5, 10, 20, $\infty$} and min_sample {0, 1, 2, 4, 8, 16, 20}.
 AFT model in XGBoost
-The model was implemented using the Python package xgboost.
-Cross-validation (cv = 5) was performed to select the optimal hyperparameters, following the same grid search described in Barnwal et al.
-[11]:
+Cross-validation (cv = 5) was performed to select the optimal hyperparameters, following the same grid search described in Barnwal et al. [14]:
   • learning_rate: 0.001, 0.01, 0.1, 1.0
@@
   • aft_loss_distribution_scale: 0.5, 0.8, 1.1, 1.4, 1.7, 2.0
+Figure 6: The mean and standard deviation of test squared hinge errors for datasets with a moderate number of features. In these scenarios, MLP generally performs well.
 KNN
-The Euclidean distance metric is applied to normalized features (mean 0, standard deviation 1) to determine the nearest neighbors, implemented using the NearestNeighbors module from the sklearn.neighbors package.
+The Euclidean distance metric is applied to normalized features (mean 0, standard deviation 1) to determine the nearest neighbors.
 A cross-validation (cv = 5) is used to select the optimal value of $K$.
 The candidate values for $K$ range from 1 to $\lceil\sqrt{n}\rceil$, where $n$ is the number of train instances.
-MMIF
-MMIF is an ensemble of 100 MMITs.
-For each MMIT, two-thirds of the train dataset is randomly selected for training, while the remaining one-third serves as the out-of-bag (OOB) set for validation.
-Additionally, one-third of the original features are randomly chosen for each MMIT.
-A cross-validation (cv = 5) is used to select the optimal hyperparameters for each MMIT including max_depth {2, 5, 10, 15, 20, 25} and min_split_samples {2, 5, 10, 20, 50}.
-Each MMIT has its own OOB error, which is used to determine the aggregation rule.
-Let $\mathcal{T}_{i}$ be an MMIT with OOB error $e_{i}$.
-The weight assigned to $\mathcal{T}_{i}$ is given by $w_{i}=\frac{\frac{1}{e_{i}}}{\sum_{j=1}^{100}\frac{1}{e_{j}}}$ and the final prediction from MMIF is computed as $\sum_{i=1}^{100}w_{i}\mathcal{T}_{i}(\mathbf{x})$ where $\mathbf{x}$ is the set of features.
 MLP
@@
 The model is trained using the Adam optimizer with a fixed learning rate of 0.001.
-Figure 2: The mean and standard deviation of the log of test squared hinge errors from simulated datasets.
-The Linear model performs best when the dataset is linear.
-In nonlinear datasets, Tree-based models achieve the best performance.
-Figure 3: The mean and standard deviation of test squared hinge errors for datasets with a high number of features (due to significantly higher error, the Constant model is omitted to improve comparative visualization).
-Tree-based models generally perform exceptionally well due to their inherent feature selection mechanism.
-On the other hand, while MLP with ReLU activation function is a more generalized Linear model, it fails to outperform the Linear model.
-One reason for this is that when a dataset contains a majority of noisy features, MLP cannot effectively reduce their impact on predictions in the same way that a Linear model with L1 regularization can.
-Figure 4: The mean and standard deviation of test squared hinge errors for datasets with a moderate number of features. In these scenarios, MLP generally performs well.
+MMIF
+MMIF is an ensemble of 100 MMITs.
+For each MMIT, two-thirds of the train dataset with one-third of the original features are randomly selected for training, while the remaining instances serves as the out-of-bag (OOB) set for validation.
+A cross-validation (cv = 5) is used to select the optimal hyperparameters for each MMIT including max_depth {2, 5, 10, 15, 20, 25} and min_split_samples {2, 5, 10, 20, 50}.
+Each MMIT has its own OOB error, which is used to determine the aggregation rule.
+Let $\mathcal{T}_{i}$ be an MMIT with OOB error $e_{i}$.
+The weight assigned to $\mathcal{T}_{i}$ is given by $w_{i}=\frac{\frac{1}{e_{i}}}{\sum_{j=1}^{100}\frac{1}{e_{j}}}$ and the final prediction from MMIF is computed as $\hat{y}=\sum_{i=1}^{100}w_{i}\mathcal{T}_{i}(\mathbf{x})$, where $\mathbf{x}$ is the set of features.
 Performance Consistency
@@
 Table 4: The comparison of the performance and consistency of 7 models across 39 datasets.
-4 Discussion
+For example, the constant model has the worst performance (rank 7th) in 20 out of 39 datasets, model mmif has the best consistency (rank 1st) in 12 out of 39 datasets.
+3.2 Results
+For each of the 39 datasets, 7 models were implemented, with each model producing 5 hinge error values.
+Figure 4 shows the comparison on 3 highly noisy simulated datasets, where 95% of the features are noise (19 out of 20).
+Figure 5 shows the comparison on 6 slightly unstable datasets, where the feature-to-instance ratio is high (from 0.06 to 1.12).
+Figure 6 shows the comparison on 6 fairly stable datasets, where the feature-to-instance ratio is low (from 0.009 to 0.09).
+For more detailed comparison of all 39 datasets—including comparisons on hinge loss, confidence intervals, and IQR boxes—see the supplemental material here: SUPPLEMENTAL MATERIALS.
+4 Discussions and Conclusions
 Dataset Variations
-To ensure a fair and comprehensive comparison, the quality of the datasets is crucial.
-First, the quality of features is considered.
-The Constant model (or Featureless), being the simplest, is expected to perform the worst across all datasets.
-However, in 19 datasets, it is not ranked the lowest.
-Notably, in 8 of these datasets, it is ranked 4th place or higher, suggesting that the features in these datasets provide limited predictive value.
-Next, the relationship between features and targets is examined.
-Since Linear models assume a linear relationship, while MMIT assumes a highly nonlinear one, dataset linearity is assessed based on performance and consistency from both Linear and MMIT models.
-If the Linear model outperforms MMIT with greater consistency, the dataset is more likely to exhibit a linear relationship.
-However, datasets where Linear performs better but with lower consistency than MMIT are not considered.
-From that criteria, out of 39 datasets, 15 are identified as more linear than nonlinear, including autohorse, autompg, bodyfat, breasttumor, cholesterol, cleveland, elusage, fishcatch, fruity, meta, pbc, simulated.linear, sleep, and vineyard.
-Conversely, 15 datasets are more nonlinear than linear, such as autoprice, basketball, cloud, cpu, lowbwt, lymphoma.mkatayama, machine.cpu, mbagrade, pollution, pwlinear, pyrim, servo, simulated.abs, simulated.sin, and strike.
-This statistic demonstrates the diversity and quality of the datasets, ensuring a fair comparison across different scenarios.
+To ensure a comprehensive comparison, the quality of the datasets is important.
+The datasets employed have a variety of characteristics, including high- and low-dimensional data, real-world and synthetic sources, linear and nonlinear patterns, as well as low- and high-noise conditions.
+An overview of the dataset qualities is provided in Table 3.
 Comprehensive Comparison
-Let’s examine the performance and consistency rankings in Table 1.
-Starting with performance, it’s clear that MMIF outperforms other models in most cases.
+The performance and consistency rankings are in Table 4, 1st means the best and 7th means the worst.
+About the performance, it’s clear that MMIF outperforms other models in most cases.
 Linear, MMIT and MLP are comparable, while Constant, KNN, and the AFT model in XGBoost show the poorest performance.
-When it comes to consistency, MMIF stands out as the most consistent.
-This is expected, as MMIF being an ensemble model, relies on a group of weak MMITs to make the final prediction, which helps improve its stability.
+About the consistency, MMIF stands out as the most consistent.
+This is expected, as MMIF being an ensemble model, relies on a group of MMITs to make the final prediction, which helps improve its stability.
 Again, Constant, KNN, and the AFT model in XGBoost show the lowest consistency.
@@
 However, if complexity is not a primary concern, and the goal is to select the model with the best overall performance and consistency, MMIF is likely the first choice.
-Alternatively, if model complexity and training time are critical factors, MMIT or Linear are preferable, as they are less complex than MMIF, require significantly fewer training resources, and still maintain reliable performance.
+Alternatively, if model complexity and training time are critical factors, MMIT is preferable, as it’s less complex than MMIF, require significantly fewer training resources, and still maintain reliable performance.
 MMIT vs MMIF
 This is a comparison worth considering.
-MMIF was introduced with the expectation that an ensemble of MMITs would outperform a single MMIT, and the results confirm this hypothesis.
-Specifically, across 39 datasets, MMIF is found to outperform MMIT in 28 of them, which constitutes the majority.
-The reason for this is straightforward: MMIF, being an ensemble of MMITs, is able to reduce overfitting by not relying on a single MMIT.
+MMIF was proposed with the expectation that an ensemble of MMITs would outperform a single MMIT, and the results confirm this hypothesis.
+Specifically, across 39 datasets, MMIF outperformed MMIT in 28 of them.
+This makes sense: MMIF, being an ensemble of MMITs, is able to reduce overfitting by not relying on a single MMIT.
 Discussion on the AFT Model in XGBoost and potential future work
-The AFT model in XGBoost is one of the most complex models in this study, but its performance and consistency do not justify the complexity.
+The AFT model in XGBoost is the complex models in this study, but its performance and consistency do not justify the complexity.
 It has a limitation in that it can only handle non-negative target intervals.
@@
 This introduces a weakness, as the model becomes highly sensitive when the dataset contains many left-censored target intervals.
-When predictions are made on the exponential scale, if the predicted value is close to zero, the mapping back to the original scale can result in large negative values, as seen in some test cases where predictions like $-10^{6}$ were made.
+When predictions are made on the exponential scale, if the predicted value is close to zero, the mapping back to the original scale can result in large negative values, as seen in some test cases where predictions like $10^{6}$ were made.
 […]
-The experiments conducted in [11] indicate that the AFT model in XGBoost performs among the worst across their six simulated datasets when compared to the Linear model and MMIT.
-A potential direction for future work is to transform left-censored intervals into interval-censored intervals by replacing all lower bounds of $\infty$- ∞ with a finite value.
+The experiments conducted in [14] (Figure 2.b) indicate that the AFT model in XGBoost performs among the worst across their six simulated datasets when compared to the Linear and MMIT.
+A potential direction for future work is to transform left-censored intervals into interval-censored intervals by replacing all lower bounds of -$\infty$ with a finite value.
 This adjustment aims to reduce the model’s sensitivity to left-censored intervals and mitigate the tendency to produce excessively large negative predictions.
@@
 SUPPLEMENTAL MATERIALS
-[…]
+Test Error for each model in each dataset
+The two models with the highest mean test Hinge Squared Error have been omitted for clearer visualization.
+Ranking table of performance and consistency of each model
 constant knn linear mmit mmif aft_xgb mlp